OpenStudy (anonymous):

The graph of a line has a slope of 2 and passes through the point (–6, –5). What is the equation of the line in point-slope form? (Points : 4) y – 6 = 2(x – 5) y + 6 = 2(x + 5) y – 5 = 2(x – 6) y + 5 = 2(x + 6)

OpenStudy (anonymous):

Use the general equation y=mx+b, then substitute in variables x and y from the point to find b.

OpenStudy (anonymous):

i tried that

OpenStudy (zpupster):

or here maybe y-y1= m(x-x1) just plug in the values y1=-5 x1=-6 m=2 be careful of the signs y-(-5) = 2(x-(-6)) or????

OpenStudy (anonymous):

and i just ended up confusing myself